solve the differential equation 2d^2y/dx^2+5dy/dx-12y=0
1
Expert's answer
2019-05-13T09:39:06-0400


Characteristic equation for given equation is

2t2 + 5t - 12=0

D = 52-4*2*(-12) = 121 = 112

roots t1= (-5+11)/(2*2) = 6/4 = 1.5, t2 = (-5-11)/(2*2) = -4

Roots of characteristic equation are real and distinct, then general solution has the form

C1*e1.5x + C2*e-4x , C1 and C2 are real constants.

Answer: C1*e1.5x +C2*e-4x, C1 and C2 are real constants.